Toner Cartridges

The office has, at points, ordered two kinds of cartridges. Both can be recycled. Currently, Nashua cartridges are easier to deal with since they are one of the standard Office Depot items. For completeness, information about both is listed below.

HP Cartridges The HP toner cartridges should not be thrown away. They should be recycled. To recycle them, put them in a box and slap a UPS label on it. The UPS label can be found in one of three ways:

1) Look in the manual inside the box. There is one there. 2) If there is no manual inside the box, then look in the office drawer (where the machine room keycard is kept) and take one from there. 3) If there are none there or you can’t find them, then call HP at 1-800-752-0900 Ext. 1872 and talk to the nice lady there. She will send you some labels.

After you have slapped the label on, call UPS and ask them to come pick up. Or check with the Campus Activities Complex. They might be able to work it into their UPS stuff.

Nashua Cartridges The Nashua toner cartridges should not be thrown away. When one is used up, remove it from the printer, with out spilling and making a mess, and place it in the bag that comes with the cartridge. Seal the bag shut. Put a new cartridge in the printer and put the old cartridge in the box so that it can be recycled. Be sure to label the box USED or DEAD so that people can tell when it’s time to reorder toner cartridges. The toner cartridges can be returned in sets of four by calling 1-800-333-3439. They will tell you what to do to return the cartridges.

New toner cartridges can be ordered through Office Depot. The item is on the check-off order form under Toner Cartridges-Nashua, 92291A Recycled. Get the requisition form approved by the right people and send it in.

Recycling

There are four blue bins at the back of the office devoted to (non-soda can) recyclables. Items without a deposit go in the bins with the recycling logo on them. Items with a bottle/can deposit go in the bins with the dollar sign ($) on them. Non-deposit items may be recycled by leaving them behind Bexley dormitory (opposite from Mass Ave) on Tuesday nights. Make sure you put them in the special bins near the curbside (take the stuff out of the bags).

The latest guidelines from abbe:

Jet-A-Way will no longer accept any recyclables left for pickups in plastic bags. All items need to be placed in the bins at curbside. Jet-A-Way has generously agreed to supply each house with 2-3 very large plastic barrels which will be kept near the pickup area. This way, if you plan to transport your recyclables to the pickup site in a plastic bag, you can empty the contents of the bag into the barrel.

Glass (clear and colored), aluminum, tin and plastics (#1 & #2 only) can be mixed together. These items need to be rinsed first.

Newspaper and white paper can be mixed and put either in brown paper bags or wrapped in twine.

The empty deposit soda cans should be sorted by maker (Coca Cola or whatever) into cardboard flats, and then taken down to La Verde’s. A set of five flats should be six dollars of deposit refund. Leave the deposit in the CokeComm can in the ’fridge. The current CokeComm is rewarding people who do this with a free can of soda for every five flats.

Empty deposit bottles, in the blue bins with the dollar sign ($), should also be taken to La Verde’s when the bins get full.

General Office Cleanliness

Office Cleaning Hour Friday afternoons at 4pm is Office Cleaning hour. During this time, people in the office should pickup and clean up around the office. If there are any dirty keyboards, please make sure to clean them too.

Recycling When you are done with your can of soda, rinse it out and put it in the cardboard soda flats at the back of the office. When you are done with your food, throw the wrappings away. Remember, do not throw out anything dangerously sharp into the trash.

Disasters In case of major disaster, there is a SIPB bucket, sponges, and cleanser on the refridgerator at the back of the room.

The Fridge

Officially, anything not maintained by CokeComm may be thrown out of the fridge. Unofficially, anything not named and dated may be thrown out, especially if it starts smelling bad (at which time it should definitely be thrown out).

The last fridge cleaning occurred in February of 1994, at which time the inside and door front of the fridge were thoroughly scrubbed with bleach-based cleanser. If you clean the fridge out, do not forget to empty the water-collection tray located beneath it. Feel free to use the gloves, sponges, bucket and cleanser located on the fridge.

Hostnames in the Office

This section was last updated on May 30th, 1998.

Hostname  IP Address Company/Model/Operating System  Role
x15-cruise-basselope 18.70.0.149  Sun/ULTRA-1/SunOS  Hotseat
snorklewacker  18.70.0.211  Dell/Pentium II/NetBSD  Spare for anti-snork
oliver  18.70.0.212  SGI/O2/IRIX  O2 testing
mary-kay-commandos 18.70.0.150  Sun/SPARC-5/Solaris  Hotseat
steve-dallas  18.70.0.213  SGI/INDY/IRIX  Hotseat
bobbi-harlow  18.70.0.62  Sun/SPARC-4/Solaris
opus  18.70.0.252  SGI/INDY/IRIX
cutter-john  18.70.0.158  PC/Pentium/Linux  Linux builds
portnoy  18.70.0.218  Sun/SPARC-5/Solaris  Backups
lola-granola  18.70.0.26  Dell/i486/NetBSD  Scanner
bart-savagewood  18.70.0.56 Sun/SPARC-CLASSIC/SOLARIS
planet-zorp  18.70.0.60  Dell/Pentium/NetBSD-1.2
yaz-pistachio  18.70.0.225 DEC/DECSTATION-5000.25/ULTRIX  Ultrix
milquetoast  18.70.0.234 Apple/PowerPC/MacOS (plus others)  Mac for users
zygorthian-space-raiders 18.70.0.61  PC/Pentium/NetBSD
bill-the-cat 18.70.0.215 Sun/SPARC-CLASSIC/Solaris
hodge-podge  18.70.0.217 DEC/DECSTATION-3100/Ultrix
pickled-herring  18.70.0.23  IBM/RS6000/AIX  AIX
w-a-thornhump-III  18.70.0.58  HP/?/HPUX  HPUX
milo  18.70.0.216NeXT/NeXTCube/NeXTstep or NetBSD  NeXT
binkley  18.70.0.214 DEC/VAXSTATION-3100/Ultrix  www load monitor
quiche-lorraine  18.70.0.227  PC/i486/Linux
lying-naked-in-the-periwinkle 18.70.0.24 DEC/DECSTATION-5000.25/Ultrix  cview

Machines It’s very important that machine hardware be kept in a proper state. This means that all connectors should be screwed into the machine where possible. The cables should also be tucked out of the way so that they don’t get accidentally tugged. Power cables should be connected to a power bar mounted on the underside of the table or desk. Ethernet cables should be connected to the nearest ethernet transceiver, also mounted on the bottom of a desk. All slack cable should be coiled up and cable-tied to keep it off the floor, so that the floor can be clean. Finally, laptop drop cables should be kept nicely coiled, and clipped and